OpenStudy (anonymous):

Provide the solution to the system. 3x – 6y = 12 9x + 2y = –24

OpenStudy (ghazi):

multiply equation one that is 3x-6y=12 by 3 and subtract it from next equation ..you'll get y

OpenStudy (anonymous):

Divide first equation by 3: x-2y=4 and rearraneg to get x=2y+4 Substitute into second: 9(2y+4)+2y=-24 Expand: 18y +36 +2y = -24 Simplify: 20y = -60 Divide by 20: y=-3. Substitute y=-3 into one of the equations, say the 1st: 3x+18=12 -> 3x=-6 ->x=-2

OpenStudy (anonymous):

@ghazi i got -3/4?

OpenStudy (ghazi):

y=3

OpenStudy (anonymous):

If you are unsure of your answer, find the corresponding x value for your y value using one of the equations. Then substitute both into the other equation to make sure it holds.

OpenStudy (ghazi):

you should get y= 3 and x=10

OpenStudy (anonymous):

Okay thank you i figured out my problem.

OpenStudy (ghazi):

hope you got the answer..it's x=-2, y=-3